y= 5/6x + 1/6
y= 7/6x + 11/6

Answer by richwmiller(17219)

since they both equal y set them equal to each other
5/6x+1/6=7/6x+11/6
subtract 1/6 from both sides
5/6x=7/6x+10/6
multiply by 6
5x=7x+10
subtract 5x
0=7x+10
subtract 10
-10=7x
divide by 7
-10/7=x
substitute in either of the original equations
y=7/6x+11/6
y=7/6*-10/7*x+11/6
y=-10/6+11/6
multiply by 6
6y=-10+11
6y=1
y=1/6
we could check by plugging the x and y into the other equation